# Beyond directed hypergraphs: heterogeneous hypergraphs and spectral centralities

### v1

@@ Abstract
The study of hypergraphs has received a lot of attention over the past few years, however up until recently there has been no interest in systems where higher order interactions are not undirected. In this article we introduce the notion of heterogeneous hypergraphs from an algebraic point of view, which have traditional directed hypergraphs as a particular case. We furthermore analytically study the spectral centralities associated to some types of heterogeneous hypergraphs, extending previously defined eigenvector-like centrality measures to this new realm. We supplement the analytical arguments with some numerical comparisons of pairwise and higher order rankings, and we construct directed higher order networks from real data.
@@ 1 Introduction
In recent years, the study of complex systems has transcended binary relationships and has embraced the multifaceted nature of interactions prevalent in real-world scenarios [1]. Traditional graph structures, while suitable for representing pairwise relationships in networks, encounter limitations when grappling with interactions involving multiple entities simultaneously. Hypergraphs [2], an extension of conventional graphs, provide a more expressive framework by accommodating hyperedges that link multiple vertices [3, 4]. This expanded structure allows for the representation of richer, multi-entity relationships, paving the way for a more comprehensive understanding of complex systems across various disciplines.
The concept of centrality, a cornerstone in network analysis, encapsulates various metrics that quantify the relative significance of nodes within a network [5, 6, 7]. A subset of these measures (the so-called spectral centralities [8]) are of paramount importance both for the possibility of analytical treatment that they provide and for the reduced computational cost that they usually entail when compared to other measures such as the betweenness centrality. However, their applicability and adaptations to hypergraphs pose intriguing challenges and opportunities for unveiling essential features of complex systems’ structures and dynamic [9]s.
The task of translating traditional network theoretical concepts such as that of centrality measures, as well as other tools and paradigms to the undirected hypergraph realm is already a challenging endeavor, so much so that little to no attention has been paid to the case of directed hypergraphs. These structures have been known for some decades now [10], although the bulk of works with them have belonged to purely mathematical or computer science literature [11, 12]. Nevertheless, it is natural to consider them as the next step for the network science community.
In the case of standard (pairwise) networks the meaning of directed interactions is clear, but when we consider interactions pertaining to multiple individuals the notion of directedness becomes fuzzy, to say the least. The traditional approach has made a somewhat sensible choice by defining a directed hyperedge as a hyperedge where some of its nodes are the “input” and some are the “output” [10, 13]. While this definition is very useful in terms of enabling a vast collection of results, it constrains the space of possible non-undirected interactions. For this reason, we find that the term “heterogeneous hyperedge” is perhaps a more adequate label for non-undirected hyperedges, including directed hyperedges as a specific type of them.
@@ 2 Preliminaries and notation
Let us start by defining a hypergraph based on the original definition, although [2] explicitly introducing the possibility of having directed hyperedges.
Definition 2.1 (Hypergraph)
@@
Hypergraphs generalize traditional graphs (where each hyperedge contains only two nodes) by allowing hyperedges to connect multiple vertices, offering a more expressive way to model relationships and interactions involving multiple elements simultaneously [3, 4].
@@
We will be analyzing the $\mathcal{H}$-eigenvector centrality (HEC) and the $\mathcal{Z}$-eigenvector centrality (ZEC) of a hypergraph $H=(V,E)$, first defined in [9]. This is based on the spectral theory of tensors [14], in particular those associated to strongly connected, $m$-uniform hypergraphs. The discussion on strong connectivity is reserved for the coming section, while uniformity is defined as follows.
Definition 2.2 ($m$-uniform hypergraph)
@@ 2.1 Algebraic properties
Definition 2.3 (Adjacency tensor of a uniform hypergraph)
@@
Definition 2.4 (Strongly connected hypergraph [9])
@@
Clearly, the strong connectivity is completely reliant on the structure of the adjacency tensor. This will become very important in the case of directed hypergraphs.
Let us now define the so-called “tensor apply” operation [15], which is nothing but the contraction of tensor $\mathcal{T}\in\mathbb{R}^{m,N}$ with vector $\mathbf{c}\in\mathbb{R}^{N}$ producing yet another vector:
@@ 2.2 The $\mathcal{H}$-eigenvector centrality
In [9], the author builds upon the theory of tensor eigenvalues [14], to define three spectral centrality measures for uniform hypergraphs: the clique-eigenvector centrality (CEC), the $\mathcal{Z}$-eigenvector centrality (ZEC) and the $\mathcal{H}$-eigenvector centrality (HEC).
Here we will only review the latter, for several reasons. First, the CEC is equivalent to the standard eigenvector centrality of the projected hypergraph (where every hyperedge is substituted by a clique among the participant nodes), therefore missing any nonlinear information contained in the higher-order structure. Second, the ZEC is a tensor-based measure which yields a fixed-norm vector (i.e. it is not re-scalable), has no uniqueness guarantees and is computationally problematic [14, 9]. It is also not compliant with a uniformization procedure as in [16]. The HEC suffers none of these problems, as we will see. What’s more, all ideas discussed in this manuscript also apply to the CEC and ZEC cases without extra effort.
@@
This eigenproblem can be cast in a network science context as a centrality measure [9], bearing in mind that it should be applied to the transposed version of the hypergraph’s associated tensor (whose meaning will be discussed later on).
Definition 2.5 (HEC of a hypergraph [9])
Let $H$ be the a $m$-uniform hypergraph, with an associated tensor $\mathcal{T}=(T_{i_{1}\dots i_{m}})$. Its $\mathcal{H}$-eigenvector centrality (HEC) is defined as the Perron-like (unique, positive) $\mathcal{H}$-eigenvector $\bf{c}$ of $(\mathcal{T})^{t}$, i.e.
$\lambda\mathbf{c}^{[m-1]}=(\mathcal{T})^{t}\mathbf{c}\dots\mathbf{c}\quad\Rightarrow\quad\lambda c_{i_{1}}^{m}=\sum_{i_{2}\dots i_{m}=1}(T_{i_{1}\dots i_{m}})^{t}\,c_{i_{2}}\dots c_{i_{k}},\quad\mathbf{c}>0,\;\lambda=\rho((\mathcal{T})^{t}).$ (II.4)
@@ 3 Heterogeneous hypergraph zoology
In this section we want to begin our quest to generalize hypergraphs to include some notion of directedness. As we advanced in the introduction, so far in the literature the notion of directed hypergraphs has been restricted to a special type of higher order interaction, especially amenable to a set-theoretic description [10]. Our first task is, therefore, breaking free from this constraint and understand how we can build none and directed interactions from other points of view. What’s more, even within the set theoretic framework we will see that it is easy to find new types of structures. The wide variety of hypergraph types requires a new adjective to describe them all, beyond traditionally directed hypergraphs, which is why we chose to call them heterogeneous hypergraphs.
@@ 3.1 Algebra, topology and set theory
Let start from the very beginning: upon opening any textbook or scientific publication on graph theory, one is most likely going to find the definition of a graph to be $G=(V,E)$, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges [6, 7]. The distinction between directed and and directed edges lies, in this context, on whether the edges $e\in E$ are ordered or unordered pairs of nodes. Notice that this is a set theoretic point of view: a graph can also be understood from a topological point of view (nodes and connections between them) or from an algebraic point of view (e.g. adjacency matrix), see Figure 1.
@@
Starting from the set-theoretic viewpoint, the simplest idea is allowing each hyperedge to have multiple “input” nodes, and multiple “output” nodes. In [10] this perspective is fleshed out: a directed hyperedge is a tuple $E_{k}=(H(E_{k}),T(E_{k}))$ where $H(E_{k})\subseteq V$ is the “head” of the hyperedge, while $T(E_{k})\subseteq V$ is the “tail” of the hyperedge, and such that [^2] According to [10] either the head or the tail set could be empty although we will assume from here on that they aren’t. $H(E_{k})\cap T(E_{k})=\emptyset$. We will elaborate on this type of hypergraph structures on 4.2. See Figure 2 for an example.
@@
This situation is reminiscent of the division between simplicial complexes and hypergraphs: the former is a more restrictive type than the latter, however if we are interested in a problem where its topological features and tools (e.g. homology) can be exploited [1], then it is advantageous to discuss it on its own. Here we can see that if the directedness is present in a binary fashion (input-output), these structures are worth considering, but we should keep in mind that there are other possible ways of non-undirectedness in the interactions of a hypergraph.
@@ 3.1.1 Applications and examples of heterogeneous interaction systems
The possibility of studying systems with heterogeneous interactions in the abovementioned sense seems so far like a mathematical pastime, bearing no connection to any real system, however that is far from reality. In fact, the original paper on directed hypergraphs [10] already describes some use cases, and the need to describe heterogeneous interactions was already recognized more than a decade ago [17].
Here we briefly review some of the applications described in both references and we add some more which we think are of relevance.
  • Biochemical reactions. Online, open databases such as [18] or [19] can be mined in order to construct different directed hypergraphs where edges represent chemical reactions or metabolical pathways between metabolites, respectively [17]. We performed this mining task, and the former example is later showcased in Subsection 4.2.
@@
  • Urban transit. In [10] it is argued that there are many mobility networks which can be abstracted as directed hypergraphs, where some nodes are stops within transportation lines (e.g. metro), and thus connected via pairwise edges representing said lines, while other nodes serve as the actual transit stops, which individuals need to traverse by walking to change lines [20]. These are represented as directed hyperedges.
@@
    A more restrictive, though more tractable version of this is the case where all paths are of the same length $k$. This is then related to the $k$-step eigenvector centrality originally put forward in [21] in the $k=2$, undirected case, which we will extend in Subsection 4.3.
@@
  • Propositional logic and relational databases. In [10] different such hypergraphs are considered, with the most prominent type abstracting relational databases, where nodes represent propositions, with directed hyperedges linking a series of causes to their consequences. Notice that here the projection of the hypergraph to its pairwise constituents is losing key information (one needs all propositions to be fulfilled at once in order to prove their consequence).
@@ 3.2 The transposition operation
Definition 3.1 (Transposition of an adjacency tensor)
Let $\mathcal{T}\in\mathbb{R}^{[m,N]}$ be an adjacency tensor, $\sigma$ a permutation of $m$ elements. The transposition of the tensor is defined, component by component, $(T_{i_{1}...i_{m}})^{t}=T_{\sigma(i_{1}...i_{m})},$ where $\sigma(i_{1}...i_{m})$ is the permutation of the indices $\{i_{1},...,i_{m}\}$.
@@
Definition 3.2 (Transposed hypergraph)
@@ 4.1 Cyclical hypergraphs
Definition 4.1 (Cyclical hyperedge)
@@
[…]
@@ 4.2 Directed hyperedges
Definition 4.2 ($\mathcal{B}$- and $\mathcal{F}$-hyperedges [10])
@@
It is common to refer to $h(e)$ as the “head” of the hyperedge, and to $t(e)$ as its “tail” [10]. Notice that any hyperedge $e=(t(e),h(e))$ not belonging to either of them (i.e. $|t(e)|,|h(e)|\neq 1$) can be converted into a $\mathcal{B}$-hyperedge and a $\mathcal{F}$-hyperedge by splitting it placing a node between them (although from the point of view of centrality this is not an appropriate operation, for it creates new nodes with their own centrality). A hypergraph consisting of just $\mathcal{B}$-hyperedges ($\mathcal{F}$-hyperedges) is called a $\mathcal{B}$-hypergraph ($\mathcal{F}$-hypergraph).
Directed hypergraphs can be encoded as a series of adjacency tensors whose components satisfy certain symmetry constrains (this has already been proposed by [13]). Each directed hyperedge contributes to the corresponding adjacency tensor as follows:
@@
Algebraicly, the transposition of such a tensor corresponds to the index transposition
@@
for each tensor $T^{(2)},T^{(3)},...,T^{(M)}$. The following result is straightforward.
Proposition 4.3 (Transposed directed hypergraph)

### v2

@@ Abstract
The study of hypergraphs has received a lot of attention over the past few years, however up until recently there has been no interest in systems where higher order interactions are not undirected. In this article we introduce the notion of heterogeneous hypergraphs from an algebraic point of view, which have traditional directed hypergraphs as a particular case. We furthermore analytically study the spectral centralities associated to some types of heterogeneous hypergraphs, extending previously defined eigenvector-like centrality measures to this new realm. We supplement the analytical arguments with some numerical comparisons of pairwise and higher order rankings, and we construct directed higher order networks from real data, which we then use for discussion and analysis.
@@ 1 Introduction
In recent years, the study of complex systems has transcended binary relationships and has embraced the multifaceted nature of interactions prevalent in real-world scenarios [battiston2021physics]. Traditional graph structures, while suitable for representing pairwise relationships in networks, encounter limitations when grappling with interactions involving multiple entities simultaneously. Hypergraphs [berge1984hypergraphs], an extension of conventional graphs, provide a more expressive framework by accommodating hyperedges that link multiple vertices [battiston2020networks, boccaletti2023structure]. This expanded structure allows for the representation of richer, multi-entity relationships, paving the way for a more comprehensive understanding of complex systems across various disciplines.
The concept of centrality, a cornerstone in network analysis, encapsulates various metrics that quantify the relative significance of nodes within a network [boccaletti2006complex, estrada2012structure, newman2010networks]. A subset of these measures (the so-called spectral centralities [vigna2016spectral]) are of paramount importance both for the possibility of analytical treatment that they provide and for the reduced computational cost that they usually entail when compared to other measures such as the betweenness centrality. However, their applicability and adaptations to hypergraphs pose intriguing challenges and opportunities for unveiling essential features of complex systems’ structures and dynamics [benson2019three].
The task of translating traditional network theoretical concepts such as that of centrality measures, as well as other tools and paradigms to the undirected hypergraph realm is already a challenging endeavor, so much so that little to no attention has been paid to the case of directed hypergraphs. These structures have been known for some decades now [gallo1993directed], although the bulk of works with them have belonged to purely mathematical or computer science literature [xie2016spectral, ausiello2017directed]. Nevertheless, it is natural to consider them as the next step for the network science community.
In the case of standard (pairwise) networks the meaning of directed interactions is clear, but when we consider interactions pertaining to multiple individuals the notion of directedness becomes fuzzy, to say the least. The traditional approach has made a somewhat sensible choice by defining a directed hyperedge as a hyperedge where some of its nodes are the “input” and some are the “output” [gallo1993directed, gallo2022synchronization]. While this definition is very useful in terms of enabling a vast collection of results, it constrains the space of possible non-undirected interactions. For this reason, we find that the term “heterogeneous hyperedge” is perhaps a more adequate label for non-undirected hyperedges, including directed hyperedges as a specific type of them.
@@ 2 Preliminaries and notation
Let us start by defining a hypergraph based on the original definition, although [berge1984hypergraphs] explicitly introducing the possibility of having directed hyperedges.
Definition 2.1 (Hypergraph).
@@
Hypergraphs generalize traditional graphs (where each hyperedge contains only two nodes) by allowing hyperedges to connect multiple vertices, offering a more expressive way to model relationships and interactions involving multiple elements simultaneously [battiston2020networks, boccaletti2023structure].
@@
We will be analyzing the $\mathcal{H}$-eigenvector centrality (HEC) and the $\mathcal{Z}$-eigenvector centrality (ZEC) of a hypergraph $H=(V,E)$, first defined in [benson2019three]. This is based on the spectral theory of tensors [qi2017tensor], in particular those associated to strongly connected, $m$-uniform hypergraphs. The discussion on strong connectivity is reserved for the coming section, while uniformity is defined as follows.
Definition 2.2 ($m$-uniform hypergraph).
@@ 2.1 Algebraic properties
Definition 2.3 (Adjacency tensor of a uniform hypergraph).
@@
Definition 2.4 (Strongly connected hypergraph [benson2019three]).
@@
Clearly, the strong connectivity is completely reliant on the structure of the adjacency tensor. This will become very important in the case of directed hypergraphs. Note that, even when a hyperedge is not undirected, this definition still makes sense, as it yields a weight for the relation between two nodes, summing others involved in the larger interaction.
Let us now define the so-called “tensor apply” operation [benson2019computing], which is nothing but the contraction of tensor $\mathcal{T}\in\mathbb{R}^{m,N}$ with vector $\mathbf{c}\in\mathbb{R}^{N}$ producing yet another vector:
@@ 2.2 The $\mathcal{H}$-eigenvector centrality
In [benson2019three], the author builds upon the theory of tensor eigenvalues [qi2017tensor], to define three spectral centrality measures for uniform hypergraphs: the clique-eigenvector centrality (CEC), the $\mathcal{Z}$-eigenvector centrality (ZEC) and the $\mathcal{H}$-eigenvector centrality (HEC).
Here we will only review the latter, for several reasons. First, the CEC is equivalent to the standard eigenvector centrality of the projected hypergraph (where every hyperedge is substituted by a clique among the participant nodes), therefore missing any nonlinear information contained in the higher-order structure. Second, the ZEC is a tensor-based measure which yields a fixed-norm vector (i.e. it is not re-scalable), has no uniqueness guarantees and is computationally problematic [qi2017tensor, benson2019three]. It is also not compliant with a uniformization procedure as in [contrerasaso2023uplifting]. The HEC suffers none of these problems, as we will see. What’s more, all ideas discussed in this manuscript also apply to the CEC and ZEC cases without extra effort.
@@
This eigenproblem can be cast in a network science context as a centrality measure [benson2019three], bearing in mind that it should be applied to the transposed version of the hypergraph’s associated tensor, $\mathcal{T}^{t}$ (whose meaning and definition will be discussed in the next Section).
Definition 2.5 (HEC of a hypergraph [benson2019three]).
Let $H$ be the a $m$-uniform hypergraph, with an associated tensor $\mathcal{T}=(T_{i_{1}\dots i_{m}})\in\mathbb{R}^{[m,N]}$. Its $\mathcal{H}$-eigenvector centrality (HEC) is defined as the Perron-like (unique, positive) $\mathcal{H}$-eigenvector $\bf{c}$ of $\mathcal{T}^{t}$, i.e.
$\lambda\mathbf{c}^{[m-1]}=(\mathcal{T})^{t}\mathbf{c}\dots\mathbf{c}\quad\Rightarrow\quad\lambda c_{i_{1}}^{m}=\sum_{i_{2}\dots i_{m}=1}(T_{i_{1}\dots i_{m}})^{t}\,c_{i_{2}}\dots c_{i_{k}},\quad\mathbf{c}>0,\;\lambda=r(\mathcal{T}^{t}),$ (II.4)
where $r(\mathcal{T}^{t})$ denotes the spectral radius of $\mathcal{T}^{t}$.
@@ 3 Heterogeneous hypergraph zoology
In this Section we want to begin our quest to generalize hypergraphs, to include some notion of directedness. As we advanced in the introduction, so far in the literature the notion of directed hypergraphs has been restricted to a special type of higher order interaction, especially amenable to a set-theoretic description [gallo1993directed]. Our first task is, therefore, breaking free from this constraint and understanding how we can build directed interactions from other points of view. What’s more, even within the set theoretic framework we will see that it is easy to find new types of structures. The wide variety of hypergraph types requires a new adjective to describe them all, beyond traditionally directed hypergraphs, which is why we chose to call them heterogeneous hypergraphs.
@@ 3.1 Algebra, topology and set theory
Let us start from the very beginning: upon opening any textbook or scientific publication on graph theory, one is most likely going to find the definition of a graph to be $G=(V,E)$, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges [estrada2012structure, newman2010networks]. The distinction between directed and and directed edges lies, in this context, on whether the edges $e\in E$ are ordered or unordered pairs of nodes. Notice that this is a set theoretic point of view: a graph can also be understood from a topological point of view (nodes and connections between them) or from an algebraic point of view (e.g. adjacency matrix), see Figure 1.
@@
Starting from the set-theoretic viewpoint, the simplest idea is allowing each hyperedge to have multiple “input” nodes, and multiple “output” nodes. In [gallo1993directed] this perspective is fleshed out: a directed hyperedge [^2] It should be noted that in some works in the literature [jost2019hypergraph, mulas2021spectral] these hyperedges are called oriented instead of directed. is a tuple $E_{k}=(H(E_{k}),T(E_{k}))$ where $H(E_{k})\subseteq V$ is the “head” of the hyperedge, while $T(E_{k})\subseteq V$ is the “tail” of the hyperedge, and such that [^3] According to [gallo1993directed], either the head or the tail set could be empty although we will assume from here on that they aren’t. $H(E_{k})\cap T(E_{k})=\emptyset$. We will elaborate on this type of hypergraph structures on 4.2. See Figure 2 for an example.
@@
This situation is reminiscent of the division between simplicial complexes and hypergraphs: the former is a more restrictive type than the latter, however if we are interested in a problem where its topological features and tools (e.g. homology) can be exploited [battiston2021physics], then it is advantageous to discuss it on its own. Here we can see that if the directedness is present in a binary fashion (input-output), these structures are worth considering, but we should keep in mind that there are other possible ways of non-undirectedness in the interactions of a hypergraph.
@@ 3.1.1 Applications and examples of heterogeneous interaction systems
The possibility of studying systems with heterogeneous interactions in the abovementioned sense seems so far like a mathematical pastime, bearing no connection to any real system, however that is far from reality. In fact, the original paper on directed hypergraphs [gallo1993directed] already describes some use cases, and the need to describe heterogeneous interactions was already recognized more than a decade ago [klamt2009hypergraphs].
Here we briefly review some of the applications described in both references and we add some more which we think are of relevance:
  • Biochemical reactions. Online, open databases such as [KIDA-datasets] or [BiGG-datasets] can be mined in order to construct different directed hypergraphs where edges represent chemical reactions or metabolical pathways between metabolites, respectively [klamt2009hypergraphs]. We performed this mining task, and the former example is later showcased in Subsection 4.2.
@@
  • Urban transit. In [gallo1993directed] it is argued that there are many mobility networks which can be abstracted as directed hypergraphs, where some nodes are stops within transportation lines (e.g. metro), and thus connected via pairwise edges representing said lines, while other nodes serve as the actual transit stops, which individuals need to traverse by walking to change lines [criado2010hyperstructures]. These are represented as directed hyperedges.
@@
    A more restrictive, though more tractable version of this is the case where all paths are of the same length $k$. This is then related to the $k$-step eigenvector centrality originally put forward in [xu2023two] in the $k=2$, undirected case, which we will extend in Subsection 4.3.
@@
  • Propositional logic and relational databases. In [gallo1993directed] different such hypergraphs are considered, with the most prominent type abstracting relational databases, where nodes represent propositions, with directed hyperedges linking a series of causes to their consequences. Notice that here the projection of the hypergraph to its pairwise constituents is losing key information (one needs all propositions to be fulfilled at once in order to prove their consequence).
@@ 3.2 The transposition operation
Definition 3.1 (Transposition of an adjacency tensor).
Let $\mathcal{T}\in\mathbb{R}^{[m,N]}$ be an adjacency tensor and $\sigma$ a permutation of $m$ elements. The transposition of the tensor is defined, component by component, as $(T_{i_{1}...i_{m}})^{t}=T_{\sigma(i_{1}...i_{m})},$ where $\sigma(i_{1}...i_{m})$ is the permutation of the indices $\{i_{1},...,i_{m}\}$.
@@
Definition 3.2 (Transposed hypergraph).
@@ 4.1 Cyclical hypergraphs
Definition 4.1 (Cyclical hyperedge).
@@
$T_{\sigma(i_{1},\dots,i_{m})}=T_{\sigma^{\prime}(i_{1},\dots,i_{m})},\;\forall\sigma,\sigma^{\prime}\in\sigma_{\rm odd}\quad\text{and}\quad T_{\sigma(i_{1},\dots,i_{m})}=T_{\sigma^{\prime}(i_{1},\dots,i_{m})},\;\forall\sigma,\sigma^{\prime}\in\sigma_{\rm even}$ (IV.5)
@@ 4.2 Directed hyperedges
Definition 4.2 ($\mathcal{B}$- and $\mathcal{F}$-hyperedges [gallo1993directed]).
@@
It is common to refer to $h(e)$ as the “head” of the hyperedge, and to $t(e)$ as its “tail” [gallo1993directed]. Notice that any hyperedge $e=(t(e),h(e))$ not belonging to either of them (i.e. $|t(e)|,|h(e)|\neq 1$) can be converted into a $\mathcal{B}$-hyperedge and a $\mathcal{F}$-hyperedge by splitting it placing a node between them (although from the point of view of centrality this is not an appropriate operation, for it creates new nodes with their own centrality). A hypergraph consisting of just $\mathcal{B}$-hyperedges ($\mathcal{F}$-hyperedges) is called a $\mathcal{B}$-hypergraph ($\mathcal{F}$-hypergraph).
Directed hypergraphs can be encoded as a series of adjacency tensors whose components satisfy certain symmetry constrains (this has already been proposed by [gallo2022synchronization]). Each directed hyperedge contributes to the corresponding adjacency tensor as follows:
@@
From an algebraic point of view, the transposition of such a tensor corresponds to the index transposition
@@
for each tensor $T^{(2)},T^{(3)},...,T^{(M)}$. The following result is then straightforward.
Proposition 4.3 (Transposed directed hypergraph).